# for Dynamic Neural Radiance Fields

### v3

@@ Abstract
The accurate reconstruction of dynamic scenes with neural radiance fields is significantly dependent on the estimation of camera poses. Widely used structure-from-motion pipelines encounter difficulties in accurately tracking the camera trajectory when faced with separate dynamics of the scene content and the camera movement. To address this challenge, we propose Dynamic Motion-Aware Fast and Robust Camera Localization for Dynamic Neural Radiance Fields (DynaMoN). DynaMoN utilizes semantic segmentation and generic motion masks to handle dynamic content for initial camera pose estimation and statics-focused ray sampling for fast and accurate novel-view synthesis. Our novel iterative learning scheme switches between training the NeRF and updating the pose parameters for an improved reconstruction and trajectory estimation quality. The proposed pipeline shows significant acceleration of the training process. We extensively evaluate our approach on two real-world dynamic datasets, the TUM RGB-D dataset and the BONN RGB-D Dynamic dataset. DynaMoN improves over the state-of-the-art both in terms of reconstruction quality and trajectory accuracy. We plan to make our code public to enhance research in this area.
@@ I INTRODUCTION
Enabling novel view synthesis on dynamic scenes often requires multi-camera setups [1]. However, everyday dynamic scenes are often captured by one single moving camera, restricting the field of view [1]. In Simultaneous Localization and Mapping (SLAM) pipelines, multiple scene representations can be used. Commonly used explicit voxel and surfel representations are complemented by implicit truncated signed distance fields (TSDF) or neural radiance fields (NeRF). The latter allows the rendering of photorealistic novel views from new, reasonable camera positions [2, 3]. While many are restricted to static scenes, dynamic NeRF [4, 5, 6, 7] allows novel view synthesis for dynamic scenes.
@@
To enable novel view synthesis using NeRF, accurate camera poses, usually retrieved via COLMAP [8] or the Apple ARKit, are essential. The structure-from-motion (SfM)-based approach often demands hours of computation to estimate reasonable camera poses and is challenged by large-scale dynamic scenes. To overcome the limitation of slow SfM approaches, recent works combined SLAM and NeRF [9, 10, 11, 2]. SLAM provides faster results for the camera trajectory than classic SfM. However, by nature, scenes are more often dynamic than static. Existing NeRF-SLAM [9, 10, 11, 2] methods assume a static scene which poses difficulties when dynamic scene content is present as shown in Fig 1.  [^1] This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
@@
Figure 1: NeRF-SLAM approaches rely on static scenes (left). Approaches such as InstantNGP [3] can be used to mask out dynamic scene content [2, 11] (left). Still, when using the masking, the scene representation lacks quality (left). Fully dynamic NeRF provides an implicit 4D (3D+time) representation; however, it relies on offline SfM, which can suffer in the presence of considerable motion (center). We consider dynamics via our motion-based segmentation (MS) and semantic segmentation (SS) already in the camera localization (right) and further during novel view synthesis. This enables a more robust camera tracking and novel view synthesis with a higher quality (ours, right).
In this paper, we present DynaMoN, a motion-aware camera localization and visualization approach that can handle highly dynamic scenes in camera pose retrieval and dynamic scene representation. We utilize semantic segmentation and generic motion masks during our camera localization process to enable robust tracking of the camera path in a dynamic environment. These masks are further utilized for ray sampling to train an efficient dynamic NeRF, based on HexPlane [4], that refines the initial camera poses, leading to improved reconstruction quality in scenes with strong motions.
@@
We conduct experiments on two challenging datasets, namely the dynamic subset of the TUM RGB-D [12] and the BONN RGB-D Dynamic [13] dataset. The considered masking in DynaMoN allows us to outperform existing approaches in terms of the absolute trajectory error and novel view synthesis results on these datasets.
@@ II RELATED WORK
As DynaMoN is a combination of dynamic NeRF with fast and robust motion-aware camera localization, we group related work in topics for camera localization and their employment for novel view synthesis and neural representation for dynamic scenes.
@@ II-A Camera Localization and Scene Dynamics
SfM and SLAM [14, 15, 16, 17] are arguably the two most common approaches for robust camera localization. More specifically, SLAM often targets real-time localization of a camera coupled with mapping. In addition to visual information, some approaches can utilize complementary sensory signals such as depth images and IMU data to further improve the accuracy. A common branch of work employs ORB descriptors to build a 3D map and track the camera [14, 17, 15].
In addition to classical or partially learnable SLAM approaches, Teed and Deng [16] introduce an end-to-end learnable method. Their approach, termed DROID-SLAM, consists of a dense bundle adjustment layer that enables recurrent iterative updates of camera poses and pixel-wise depth.
Typical camera localization approaches rely on a static scene assumption. However, natural scenes are often dynamic [18, 19, 20, 21]. To address this, Yu et al. [18] introduce a dynamic SLAM approach building upon ORB-SLAM2 [17]. Their DS-SLAM integrates a semantic segmentation network and a moving consistency check in the tracking process. Rünz et al. [22] enrich RGB-D SLAM with object-awareness and semantic segmentation capabilities, while DynaSLAM [23] combines geometric and deep learning-based segmentation to mask dynamic objects in an attempt to ignore them for mapping. Others split the scene into dynamic and static parts. Dai et al. [20] build upon ORB-SLAM2 [17] and introduce a semantic segmentation using point correlation for this split and Ye et al. [21] divide the scene representation in a static flow field and a dynamic flow field. Yu et al. [24] combine semantic segmentation with optical flow to remove dynamic keypoints achieving a higher robustness compared to ORB-SLAM2.
Following this idea, we also split the scene into static parts and dynamic content using predicted segmentation masks.
II-B Camera Localization and NeRF
With the retrieved camera poses, a captured environment can be represented in 3D. Aside from explicit maps, an implicit scene representation can be established using a multilayer perceptron (MLP) [25]. Mildenhall et al. [25] introduce the use of an MLP as a neural representation for novel view synthesis, denoted as NeRF. The implicit representation is regressed by a photometric comparison of rendered views with input images. This allows the synthesis of photorealistic images from novel viewpoints at the cost of the requirement for accurate camera poses during training, typically captured by computationally intensive SfM with COLMAP [8]. An efficient alternative to SfM is the use of SLAM to retrieve camera poses.
In this regard, Sucar et al. [26] propose the usage of an MLP as scene representation in a SLAM framework and Zhu et al. [9] arrange multiple NeRFs in a hierarchical feature grid to learn both camera pose and scene representation. NICER-SLAM [10] introduces a similar approach for RGB input with estimated depth. Others purely rely on RGB and combine SLAM and NeRF [2, 11] utilizing hash-based InstantNGP [3] for scene representation. NeRF-SLAM [2] adapts the initial Droid-SLAM [16] framework and Orbeez-SLAM [11] leverages visual odometry [17].
To incorporate efficient pose estimation, we also rely on design choices of recent SLAM frameworks; however, we combine them with computationally efficient proposals [4] for neural radiance fields.
@@ II-C Neural Representation of Dynamic Scenes
Our world is inherently dynamic which violates commonly used static scene assumptions [9, 10, 26, 11, 2, 3, 25]. To represent dynamics in a NeRF [27, 28], the use of grid structures became popular [4, 5]. HexPlane [4] utilizes a 4D space-time grid divided into six feature planes. The feature vector is a 4D point in space-time projected onto each feature plane. Similarly, Tensor4d [5] builds upon a 4D tensor field using time-aware volumes projected onto nine 2D planes. TiNeuVox [7] represents dynamic scenes using time-aware voxel features to enable faster training.
Optimizing for both camera poses and view synthesis can improve results in static scenes [29, 30, 31, 32, 33]. Liu et al. [34] therefore introduce such an optimization for a dynamic environment. To model this, they use one static and one dynamic NeRF.
We, in contrast, split the problem into the differentiation of static and dynamic content already used for camera pose estimation. The dynamic environment is then represented by a dynamic, 4D NeRF calculated upon the camera poses from the static scene components, while further refining these poses based on the static parts of the images.
@@ III METHOD
We consider our input to be an RGB video captured in a dynamic environment using a single monocular moving camera without pose knowledge. This leads to two objectives: first, estimating the camera pose, and second, reconstructing the scene, which may comprise some deformation. We leverage masks to separate the scene dynamics from the static background for the former and utilize a 4D spatio-temporal neural representation for the latter. This allows to optimize for novel view synthesis while refining initial camera poses.
Figure 2: DynaMoN architecture. DynaMoN retrieves consecutive RGB frames and applies semantic segmentation ($M_{ss}\rightarrow S_{S_{i}}$) and motion masks ($M_{MS}\rightarrow S_{M_{i}}$) on these input frames ($I$). Combining these masks (OR), we enable a motion-aware, fast and robust camera pose estimation. Based on the predicted camera poses $G$, we produce a time-dependent output ($\Theta$) using dynamic NeRF. During NeRF training, we utilize the previously computed masks to refine the camera poses (dashed lines). Dashed lines represent iterative updates on NeRF and pose parameters. We highlight our approach in green.
@@ III-A Robust Camera Localization in Dynamic Environments
Novel view synthesis methods commonly employ rigid SfM, often COLMAP [8], for camera pose estimation. This assumes a static scene instead of dynamic scenes.
To cope with dynamics, it is possible to add binary masks to remove non-static regions. However, the high computation time of SfM approaches remains a drawback. In our approach, we target both robustness against dynamicscenes and a more efficient solution for pose estimation for dynamic moving cameras. Our underlying architecture is based on DROID-SLAM [16]. While it incorporates elements to cope with a certain level of scene dynamics, we found it can be improved by eliminating the contribution of the non-static regions for camera tracking. We do this by generating binary dynamics masks for each frame, which are combinations of semantic ($M_{SS}$) and motion masks ($M_{MS}$).
@@
III-B Neural Dynamic Scenes
Our scene representation builds on a state-of-the-art dynamic NeRF, HexPlane [4]. Targeting efficient optimization, fast rendering and high reconstruction quality, HexPlane [4] utilizes TensoRF [36] and constructs a 4D representation with three spatial and one temporal axes to create feature planes with their pairwise combinations to be the input of their tiny MLP. The grid size of each axis is governed in a coarse-to-fine fashion along the training to improve convergence and naturally regularizes the local neighborhoods. For optimization, we follow the original NeRF ray casting and apply the pixel-wise photometric loss between the rendered RGB values $C$ and the ground truth values $\hat{C}$, supported by a total variation regularization as:
@@ III-B Neural Dynamic Scenes
TABLE I: RMSE of the translational ATE [m] on TUM RGB-D [12]. The upper rows are the slightly dynamic scenes and the lower rows are the highly dynamic scenes. The - denotes that no results were reported for these sequences.
Sequences RGB-D RGB
DVO-SLAM ORB-SLAM2 PointCorr DynaPix-D [37] Yu et al. DytanVO DeFlowSLAM DROID-SLAM OURS (MS&SS) OURS (MS)
[38] [17] [20] [37] [24] [35] [21] [16]
fr2/desk-person 0.104 0.006 0.008 - - 1.166 0.013 0.006 0.007 0.006
fr3/sitting-static 0.012 0.008 0.010 - 0.006 0.016 0.007 0.005 0.005 0.005
fr3/sitting-xyz 0.242 0.010 0.009 - - 0.260 0.015 0.009 0.010 0.009
fr3/sitting-rpy 0.176 0.025 0.023 - - 0.046 0.027 0.022 0.024 0.021
fr3/sitting-halfsphere 0.220 0.025 0.024 - 0.018 0.310 0.025 0.014 0.023 0.019
fr3/walking-static 0.752 0.408 0.011 0.007 0.009 0.021 0.007 0.012 0.007 0.014
fr3/walking-xyz 1.383 0.722 0.087 0.014 0.014 0.028 0.018 0.016 0.014 0.014
fr3/walking-rpy 1.292 0.805 0.161 0.123 0.031 0.155 0.057 0.040 0.031 0.039
fr3/walking-halfsphere 1.014 0.723 0.035 0.023 0.020 0.385 0.420 0.022 0.019 0.020
Mean/Max 0.577/1.383 0.304/0.805 0.041/0.161 - - 0.265/1.166 0.065/0.420 0.016/0.040 0.016/0.031 0.016/0.039
@@ III-C Pose Refinement and Training Strategy
We utilize eqn. (1) as the objective function for training both the NeRF and refining the camera poses. For the latter, we employ an alternating training strategy in which every 5 000 iterations, the parameters to be optimized are switched between the radiance field and the pose parameters. To improve the accuracy and robustness of the optimization, during the pose refinement phase, we constrain the sampled rays, $\mathcal{R}$, contributing to the loss to be elements of the static regions designated by the dynamics masks. Furthermore, we also use the same masks to ensure that the total variation regularization is imposed only on the static components. As the amount of non-dynamic sampled rays can vary from image to image, we scale the learning rate $\eta_{s}$ output by a learning rate scheduler by the fraction of static pixels in the image to keep the optimization steps in the same magnitude:
@@
Figure 3: RMSE of the rotational ATE [°] on the TUM RGB-D dataset (left) and the BONN RGB-D Dynamic dataset (right). The median is shown in orange and the average in dotted lines.
@@ IV EVALUATION
To evaluate our approach, we consider two aspects. First, we consider the translational and rotational absolute trajectory error (ATE) [39] using root mean square error (RMSE) in meters and degrees. Second, we analyze the rendering quality of the novel dynamic views by reporting PSNR and SSIM [40] values.
TABLE II: RMSE of the translational ATE [m] on BONN RGB-D Dynamic [13]. RGB-D results are reported by Palazzolo et al. [13]. For DROID-SLAM and ours, we calculated the results.

### v4

# Alexander Ladikos4, Nassir Navab1,5, Daniel Roth2, and Benjamin Busam1

@@ Abstract
The accurate reconstruction of dynamic scenes with neural radiance fields is significantly dependent on the estimation of camera poses. Widely used structure-from-motion pipelines encounter difficulties in accurately tracking the camera trajectory when faced with separate dynamics of the scene content and the camera movement. To address this challenge, we propose Dynamic Motion-Aware Fast and Robust Camera Localization for Dynamic Neural Radiance Fields (DynaMoN). DynaMoN utilizes semantic segmentation and generic motion masks to handle dynamic content for initial camera pose estimation and statics-focused ray sampling for fast and accurate novel-view synthesis. Our novel iterative learning scheme switches between training the NeRF and updating the pose parameters for an improved reconstruction and trajectory estimation quality. The proposed pipeline shows significant acceleration of the training process. We extensively evaluate our approach on two real-world dynamic datasets, the TUM RGB-D dataset and the BONN RGB-D Dynamic dataset. DynaMoN improves over the state-of-the-art both in terms of reconstruction quality and trajectory accuracy. We plan to make our code public to enhance research in this area. Code available: https://hannahhaensen.github.io/DynaMoN/.
Index Terms:
Localization, Mapping
@@ I INTRODUCTION
Enabling novel view synthesis (NVS) on dynamic scenes often requires multi-camera setups [1]. However, everyday dynamic scenes are often captured by one single moving camera, restricting the field of view [1]. To receive NVS or a 3D/4D scene representation, accurate camera poses are essential for such a representation. One way to retrieve camera poses is Simultaneous Localization and Mapping (SLAM). SLAM can result in multiple scene representations such as explicit voxel or surfel representations or neural radiance fields (NeRF) representations. The latter allows the rendering of photorealistic novel views from new, reasonable camera positions [2, 3]. While many are restricted to static scenes, dynamic NeRF [4, 5, 6, 7] allows NVS for dynamic scenes.
@@
To enable NVS, accurate camera poses, usually retrieved via structure-from-motion (SfM) [8] or the Apple ARKit, are essential. SfM often demands hours of computation to estimate reasonable camera poses and is challenged by large-scale dynamic scenes. To overcome the limitation of time-intensive SfM, recent works combined SLAM and NeRF [9, 10, 11, 2]. SLAM provides faster results for the camera trajectory than classic SfM. However, by nature, scenes are more often dynamic than static. Existing Radiance Field-SLAM [9, 10, 11, 2, 12, 13] methods assume a static scene which poses difficulties when dynamic scene content is present as shown in Fig 1.
@@
Figure 1: NeRF combined with SLAM usually relies on static scenes (left). Approaches such as InstantNGP [3] can be used to mask out dynamic content [2, 11] (left) resulting still in reduced quality. Fully dynamic NeRF provides an implicit 4D (3D+time) representation; however, it relies on offline SfM, which can suffer in the presence of considerable motion (center). DynaMoN considers dynamics at all stages using motion segmentation (MS) and semantic segmentation (SS) (right). This enables a more robust camera tracking and NVS with a higher quality (ours, right).
In this paper, we present DynaMoN, a motion-aware camera localization and visualization approach that can handle highly dynamic scenes.
  1. 1.
    We propose a novel combination of semantic segmentation and generic motion masks during our camera localization to enable robust tracking of the camera path in a dynamic environment for dynamic neural radiance fields.
  2. 2.
    A statics-focused ray sampling building upon our dynamic masks is introduced to enhance our dynamic NeRF representation by refining the camera poses.
Overall, (1) and (2) improve the camera poses and in turn improve the NVS, leading to improved NVS quality in scenes with strong motions.
@@
We conduct experiments on two challenging datasets, namely the dynamic subset of the TUM RGB-D [14] and the BONN RGB-D Dynamic [15] dataset. The considered masking in DynaMoN allows us to outperform existing approaches in terms of trajectory error and NVS results on these datasets.
@@ II RELATED WORK
DynaMoN combines dynamic NeRF with fast and robust motion-aware camera localization. Therefore, related work covers camera localization, camera localization for NVS, and neural representation for dynamic scenes.
@@ II-A Camera Localization and Scene Dynamics
Previous work on camera localization based on visual odometry (VO) [16], SfM [8] and SLAM [17, 18, 19, 20] extensively study architectures built for static scene assumptions. However, these methods show limited performance and often fail when they are encountered with the dynamic content commonly existent in real-world scenes.
In recent years, various classical and learning-based methods have been introduced to induce and handle the dynamic motion of the scene for robust camera pose estimation. With an assumption of prior knowledge of the objects that can independently move in the scene, DS-SLAM [21] builds upon ORB-SLAM2 [20] and integrates a semantic segmentation network to mask out motion-prone content along with a moving consistency check in the tracking process. Similarly, Dai et al. [22] extend ORB-SLAM2 [20] to handle the dynamic content with sparse point-level dynamic motion segmentation estimated by the correlation of their individual transformations across the temporal dimension. Ye et al. [23] utilize dense optical flow and leverage a learning-based approach to estimate and mask out the content that moves independently of the rigid camera motion. Building a VO pipeline for highly dynamic stereo surgical videos, Hayoz et al. [24] combine background segmentation with an end-to-end learning-based method to estimate the contribution of each pixel’s motion for camera pose optimization that is structured in a differentiable training architecture.
We build DynaMoN based on the findings of the prior work and utilize both the semantic segmentation [21] of motion-prone objects and dense learning-based motion segmentation [23] to mask out dynamic content within a SLAM architecture [19] to estimate camera pose. Furthermore, we leverage this information with our introduced static-focused ray sampling method to further fine-tune the camera poses during neural reconstruction.
II-B Camera Localization and Radiance Fields
Neural implicit methods gained popularity for scene representation. Mildenhall et al. [25] introduce NeRF employing a multi-layer perceptron (MLP) to represent a static 3D scene enabling NVS. NeRF requires precise camera poses during training, which are typically obtained through SfM pipelines like COLMAP [8] that can take long computation times for large sets of input images.
Proposing an alternative solution that can jointly optimize camera poses and scene representation, Sucar et al. [26] reformulate NeRF in a SLAM framework for ordered sets of image sequences. Zhu et al. [9] extend this by arranging multiple NeRFs in a hierarchical structure, improving the reconstruction capabilities. Following this, NICER-SLAM [10] introduces the usage of estimated depth images for more accurate geometric reconstruction. Others [2, 11] utilize hash-based InstantNGP [3] for its fast novel view rendering capabilities. NeRF-SLAM [2] adapts the Droid-SLAM [19] framework and Orbeez-SLAM [11] leverages ORB-SLAM [20].
More recent works [12, 27, 13] utilize an explicit scene representation for static scenes, namely Gaussian Splatting (GS) [28] combined with SLAM architectures for even faster convergence and novel view rendering.
Unlike the prior work, our proposed method does not carry static-scene assumptions. To this end, we utilize a two-step approach that efficiently estimates the camera poses using a dynamics-aware SLAM method and represents the scene in a dynamic NeRF architecture [4] that is extended to further refine the initially estimated camera poses.
Figure 2: DynaMoN retrieves consecutive RGB frames and applies semantic segmentation ($M_{ss}\rightarrow S_{S_{i}}$) and motion masks ($M_{MS}\rightarrow S_{M_{i}}$) on these input frames ($I$). Combining these masks (OR), we enable a motion-aware, fast and robust camera pose estimation $G$. Based on $G$, we produce a time-dependent output ($\Theta$) using dynamic NeRF. During NeRF training, we utilize the previously computed masks to refine the camera poses (dashed lines). Dashed lines represent iterative updates on NeRF and pose parameters. We highlight our approach in green.
@@ II-C Neural Representation of Dynamic Scenes
Our world is inherently dynamic which violates commonly used static scene assumptions and representations [9, 10, 26, 11, 2, 3, 29, 25]. To represent dynamics [30, 31], the use of grid structures is common [4, 5, 7]. HexPlane [4] utilizes a 4D space-time grid divided into six feature planes. The feature vector is a 4D point in space-time projected onto each feature plane. Similarly, Tensor4d [5] builds upon a 4D tensor field using time-aware volumes projected onto nine 2D planes. TiNeuVox [7] represents dynamic scenes using time-aware voxel features to enable faster training.  [32, 33] propose the usage of GS with point cloud initializations for dynamic scene representations.
Optimizing camera poses and view synthesis can improve results in static scenes [34, 35, 36, 29]. Liu et al. [37] introduce such an optimization for a dynamic NeRF using a static and a dynamic part.
We, in contrast, split the problem into the differentiation of static and dynamic content. Already at the initial camera pose estimation this split is considered. The dynamic environment is then represented by an efficient dynamic, 4D NeRF [4] utilizing the camera poses estimated from the static scene components, and further refining these poses based on the static parts of the images.
@@ III METHOD
We utilize robust and fast camera localization and an implicit scene representation, a dynamic NeRF, to enable NVS on dynamic scenes. As input, we consider an RGB video captured in a dynamic environment using a single monocular moving camera without pose knowledge. This leads to two objectives: first, estimating the camera pose, and second, reconstructing the scene, which may comprise some deformation. We leverage masks to separate the scene dynamics from the static background for the former and utilize a 4D spatio-temporal neural representation for the latter. This allows to optimize for NVS while refining initial camera poses.
@@ III-A Robust Camera Localization in Dynamic Environments
NVS methods commonly employ rigid SfM, often COLMAP [8], for camera pose estimation which assumes a static scene. To cope with dynamics, it is possible to add binary masks to remove non-static regions. However, the high computation time of SfM approaches remains a drawback. In our approach, we target both robustness against dynamic scenes and a more efficient solution for pose estimation for dynamic moving cameras. Our underlying architecture is based on DROID-SLAM [19]. While it incorporates elements to cope with a certain level of scene dynamics, we found it can be improved by eliminating the contribution of the non-static regions for camera tracking. We do this by generating binary dynamics masks for each frame, which are combinations of semantic ($M_{SS}$) and motion masks ($M_{MS}$).
@@
III-B Neural Dynamic Scene Representation
To represent the dynamic scene, we build upon a state-of-the-art dynamic NeRF, HexPlane [4]. Targeting efficient optimization, fast rendering and high reconstruction quality, HexPlane [4] utilizes TensoRF [39] and constructs a 4D representation with three spatial and one temporal axes to create feature planes. Their pairwise combinations are then the input of a tiny MLP. The grid size of each axis is governed in a coarse-to-fine fashion along the training to improve convergence and naturally regularizes the local neighborhoods. For optimization, we follow the original NeRF ray casting and apply the pixel-wise photometric loss between the rendered RGB values $C$ and the ground truth values $\hat{C}$, supported by a total variation regularization as:
@@ III-C Pose Refinement and Training Strategy
We utilize eqn. (1) as the objective function for training both the NeRF and refining the camera poses. During optimization, we employ an alternating training strategy in which every 5 000 iterations, the parameters to be optimized are switched between the radiance field and the pose parameters. To improve the accuracy and robustness of the optimization, during the pose refinement phase, we constrain the sampled rays, $\mathcal{R}$, contributing to the loss to be elements of the static regions designated by the inverse dynamics masks. Furthermore, we also use the same masks to ensure that the total variation regularization is imposed only on the static components. As the amount of non-dynamic sampled rays can vary from image to image, we scale the learning rate $\eta_{s}$ by the fraction of static pixels in the image to keep the optimization steps in the same magnitude:
@@ IV EVALUATION
To evaluate our approach, we consider the translational absolute trajectory error (ATE) [40] using root mean square error (RMSE) in meters. To reason about the orientation, we employ the Relative Rotation Accuracy (RRA) as defined in [41] after synchronizing the poses to the GT to be invariant against the alignment of coordinate frames. We calculate the median RRA over all sequences and report the percentage of poses for which the RRA lies below a specified threshold. For NVS quality, we analyze the rendering quality of the novel dynamic views by reporting PSNR and SSIM [42].
TABLE I: Percentage of orientations for which the median RRA is below the specified threshold [%] on the TUM RGB-D dataset and the Bonn RGB-D dynamic dataset.
Dataset Metric DROID-SLAM OURS (MS&SS) OURS (SS) OURS (MS)
TUM RGB-D median RRA@5 $\uparrow$ 100.0 100.0 100.0 100.0
median RRA@15 $\uparrow$ 100.0 100.0 100.0 100.0
Bonn RGB-D median RRA@5 $\uparrow$ 61.05 61.89 61.86 61.88
median RRA@15 $\uparrow$ 91.29 91.45 91.46 91.27
IV-A Datasets
To assess DynaMoN, we evaluate both the camera localization and the NVS part with state-of-the-art approaches on two dynamic datasets. In total, we evaluate our approach on 33 dynamic sequences (9 from TUM-RGB-D, and 24 from BONN RGB-D). For NVS, we follow the usual splitting using every eighth frame for testing.
TUM RGB-D - Dynamic subset [14]
. We used the nine dynamic sequences, with five sequences rated slightly dynamic and four highly dynamic. The images of the TUM RGB-D dataset have a resolution of $640\times 480$.
BONN RGB-D Dynamic [15]
The BONN RGB-D Dynamic dataset consists of 24 dynamic sequences and two static sequences with an image resolution of $640\times 480$.
TABLE II: RMSE of the translational ATE [m] on TUM RGB-D [14]. The upper rows are the slightly dynamic scenes and the lower rows are the highly dynamic scenes. The - denotes that no results were reported for these sequences.